# A torsion theoretic interpretation for sheaves of modules and Grothendieck topologies on directed categories

### v2

@@ 4. A torsion theoretic characterization of sheaves
which is exact. Restricting it to the full subcategory ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, we get a functor $\Gamma_{x}$ which is only left exact since the abelian structure of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is not inherited from that of $\mathcal{O}\operatorname{-Mod}$. Thus one can define $\mathrm{R}^{i}\Gamma_{x}(V)$, the $i$-th right derived functor of $\Gamma_{x}$, called the $i$-th sheaf cohomology group of $V$ at $x$. For details, please refer to [23, I.21.2].
@@
$\mathrm{R}^{i}\Gamma_{x}\cong\Gamma_{x}^{p}\circ(\mathrm{R}^{i+1}\mathcal{T}_{J})$
@@
It has been proved in [7, Theorem 3.16] for the special case that $J$ is the atomic Grothendieck topology and $\mathcal{O}$ is a constant structure sheaf. The proof actually works in the general situation. ∎
@@
This is essentially a combination of [13, C2.3, Lemma 2.3.3] and [13, C2.3, Corollary 2.3.4], so we only give a sketch of the proof. The functor $\phi$ induces a restriction functor
@@ 5. Grothendieck topologies on artinian EI categories
The Serre quotient on the right hand side is still complicated, so we want to find a more explicit description for some special Grothendieck topologies $J$. When $J$ is a subcategory topology (or called rigid topology), this Serre quotient is equivalent to $\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$, where $\mathcal{O}_{\mathscr{D}}$ is the restriction of $\mathcal{O}$ to a full subcategory $\mathscr{D}$ of ${\mathscr{C}}$, [^4] Since $\mathcal{O}$ is a covariant functor from $\mathcal{C}$ to the category of associative rings, its shall be restricted to ${\mathscr{D}}$ rather than ${\mathscr{D}}^{{\mathrm{op}}}$. and hence becomes much more transparent. Furthermore, it has been pointed out in [2, IV.9] that all Grothendieck topologies on a finite category are rigid, and this result also holds for artinian posets (see [17, Theorem 2.12]).
@@ 5.1. Definitions and the main result
When ${\mathscr{C}}$ is a directed category, one can construct a natural filtration on $\operatorname{Ob}({\mathscr{C}})$ as follows. Explicitly, let $O_{0}=\emptyset$, $O_{1}$ be the set of objects in $\operatorname{Ob}({\mathscr{C}})$ minimal with respect to $\leqslant$. For an ordinal $\chi$, let $O_{\chi}$ be the union of $O_{\chi-1}$ and the set of minimal elements in $\operatorname{Ob}({\mathscr{C}})\backslash O_{\chi-1}$ provided that $\chi$ is a successor ordinal, and let $O_{\chi}$ be the union of $O_{\lambda}$ for $\lambda<\chi$ if $\chi$ is a limit ordinal. In this way we obtain a chain $O_{1}\subseteq O_{2}\subseteq\cdots\subseteq\operatorname{Ob}({\mathscr{C}})$. Note that ${\mathscr{C}}$ is artinian if and only if there exists an ordinal $\zeta$ such that $O_{\zeta}=\operatorname{Ob}({\mathscr{C}})$. Dually, by taking maximal objects one gets another filtration $\{O^{\prime}_{\chi}\}$ on $\operatorname{Ob}({\mathscr{C}})$, and ${\mathscr{C}}$ is noetherian if and only if there exists an ordinal $\zeta$ such that $O^{\prime}_{\zeta}=\operatorname{Ob}({\mathscr{C}})$. For details, see [8, Subsection 3.1].
@@
Now we introduce the rigid topology and the main theorem of this section. Let ${\mathscr{C}}$ be a small skeletal category, and $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Recall from [13, C2.2] that $x\in\operatorname{Ob}({\mathscr{C}})$ is called $J$-irreducible if $J(x)=\{{\mathscr{C}}(x,-)\}$. For $y\in\operatorname{Ob}({\mathscr{C}})$, let $S_{y}$ be the sieve generated by the family of morphisms (in ${\mathscr{C}}$ rather than ${\mathscr{C}^{\mathrm{op}}}$) from $y$ to all $J$-irreducible objects. It can be explicitly described as follows
@@
A rigid Grothendieck topology $J$ is also called a subcategory topology with respect to the full subcategory consisting of $J$-irreducible objects. When $J$ is a rigid topology, $S_{y}$ is the unique minimal covering sieve in $J(y)$, so $J(y)=\{S\subseteq{\mathscr{C}}(y,-)\mid S\supseteq S_{y}\}$. For details, see [13, C2.2, Example 2.2.4(d), Remark 2.2.5, and Definition 2.2.18].
@@
Let ${\mathscr{C}}$ be a directed category. Then every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if ${\mathscr{C}}$ is a noetherian EI category.
@@ 5.2. The necessary direction
In this subsection we prove the only if direction of Theorem 5.7. Throughout this subsection let ${\mathscr{C}}$ be a directed category. As the starting point, we consider the special case that ${\mathscr{C}}$ has only one object, so we can identify ${\mathscr{C}^{\mathrm{op}}}$ with a monoid $M$, and correspondingly, a sieve on ${\mathscr{C}^{\mathrm{op}}}$ is nothing but a right ideal of $M$.
@@
A monoid $M$ has exactly two Grothendieck topologies if and only if $M$ is a group.
@@
Let ${\mathscr{C}}$ be a directed category. If every Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is rigid, then ${\mathscr{C}}$ is a noetherian EI category.
@@
We first show that ${\mathscr{C}}$ is an EI category; that is, for every $x\in\operatorname{Ob}({\mathscr{C}})$, the monoid $M_{x}={\mathscr{C}}(x,x)$ is a group. Let ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of objects $y$ with $y\geqslant x$. Clearly, ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$. By Lemma 5.10, every Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$ shall be rigid. Now let $\mathscr{E}$ be the full subcategory of ${\mathscr{D}}$ consisting of the single object $x$. Then $\mathscr{E}$ is a coideal of ${\mathscr{D}}$. By Lemma 5.11, every Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$ shall be rigid as well. By Lemma 5.9, $M_{x}^{{\mathrm{op}}}$ is a group, so is $M_{x}$. Therefore, ${\mathscr{C}}$ is an EI category.
@@ 5.3. Sufficient direction
(§)§($\S$)( § ) each $\alpha\in S_{x}$ can be written as $\alpha=\beta\gamma$ with $\beta\in S_{y}$ and $\gamma\in T$.
@@
Now we check the claim (§)§($\S$)( § ). Since $S\neq{\mathscr{C}}(x,-)$, by the consistent condition we can write $\alpha$ as a composite $\beta_{1}\gamma_{1}$ with $\gamma_{1}\in{\mathscr{C}}(x,x_{1})$ for a certain object $x_{1}\neq x$ and $\beta_{1}\in S_{x_{1}}$. We have two cases:
@@
Continuing the above procedure we obtain a sequence of objects $x<x_{1}<x_{2}<\ldots$ which must be finite since ${\mathscr{C}}$ is noetherian. Let $x_{n}$ be the last object, then either $S_{x_{n}}={\mathscr{C}}(x_{n},-)$ (and hence (§)§($\S$)( § ) holds) or $x_{n}$ is maximal with respect to $\leqslant$. For the second case, we have $\beta_{n}\in S_{x_{n}}$ is an automorphism since ${\mathscr{C}}(x_{n},x_{n})$ is assumed to be a group, so $S_{x_{n}}={\mathscr{C}}(x_{n},-)$ as well. Consequently, the claim (§)§($\S$)( § ) holds in both cases.
@@
    ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})\simeq\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}$
@@
(2): This is essentially [13, C2.2, Theorem 2.2.3]. For the convenience of the reader, we give a sketched proof. Given an $\mathcal{O}$-module $V$, its restriction to ${\mathscr{D}}$ is clearly an $\mathcal{O}_{{\mathscr{D}}}$-module. In this way we obtain an restriction functor $\textsl{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$. It has a right adjoint given by the right Kan extension described as follows: given an $\mathcal{O}_{{\mathscr{D}}}$-module $W$ and an object $x\in\operatorname{Ob}({\mathscr{C}})$,
@@
where $\iota:{\mathscr{D}}\to{\mathscr{C}}$ is the inclusion functor, and the limit is taken over the comma category $(\iota\downarrow\mathrm{const}_{x})$. When $x\in\operatorname{Ob}({\mathscr{D}})$, it is not hard to see that $(\iota\downarrow\mathrm{const}_{x})$ is a coslice category whose objects are morphisms in ${\mathscr{D}}$ ending at $x$. In this situation, the identity morphism on $x$ is the initial object of the coslice category. It follows that $(\mathrm{Ran}_{\iota}W)_{x}\cong W_{x}$ for $x\in\operatorname{Ob}({\mathscr{D}})$. Consequently, $\textsl{res}\circ\mathrm{Ran}_{\iota}$ is isomorphic to the identity functor on $\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}$. By [9, Proposition III.2.5], we conclude that
$\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}\simeq\mathcal{O}\operatorname{-Mod}/\ker(\textsl{res})=\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J)\simeq{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})$
since the kernel of res is precisely $\mathcal{T}(J)$ by (1). ∎
@@
Note that the above result holds for any rigid Grothendieck topology $J$ on any small category ${\mathscr{C}}^{{\mathrm{op}}}$ by letting ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of $J$-irreducible objects; see [13, C2.2, Theorem 2.2.3] and [13, C2.2, Example 2.2.4(d)].
@@ 6. Grothendieck topologies on certain noetherian EI categories
{generic Grothendieck topology J on $\{\text{generic Grothendieck topology $J$ on }{\mathscr{C}^{\mathrm{op}}}\}\longrightarrow\{\text{function }\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})={\mathbb{N}}\to{\mathbb{N}}\cup\{\infty\}\}.$
@@
Let $J$ be an exceptional Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ corresponded to a function $\boldsymbol{d}$, and let ${\mathscr{D}}$ be the full subcategory consisting of objects $n$ such that $d_{n}=0$ or $d_{n}=\infty$. Then ${\mathscr{D}}^{{\mathrm{op}}}$ is a $J$-dense subcategory of ${\mathscr{C}}^{{\mathrm{op}}}$ (for a definition of $J$-dense subcategories, see [13, C2.2, Definition 2.2.1]). Therefore, applying the argument in the proof of Theorem 5.16, one has
@@
and $\mathcal{O}_{{\mathscr{D}}}$ is the restriction of $\mathcal{O}$ to ${\mathscr{D}}$. The full subcategory ${\mathscr{D}}$ is an artinian category whose objects can still be parameterized by $\mathbb{N}$, and $J_{{\mathscr{D}}}$ corresponds to the sequence $(0,0,\ldots,0,\infty,\infty,\ldots)$.
@@
One can deduce the above results from Lemma 5.10. Indeed, let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Then for each $n\in\mathbb{Z}$, the full subcategory ${\mathscr{C}}_{\geqslant n}$ of objects $m\geqslant n$ is an ideal of ${\mathscr{C}}$ and an EI category of type ${\mathbb{N}}$. These results follows via applying Propositions 6.3, 6.6 and 6.8 to the restriction of $J$ to ${\mathscr{C}}_{\geqslant n}$. When ${\mathscr{C}}$ is the poset $\mathbb{Z}$, they have been proved in [17, Proposition B.31].

### v3

@@ 4. A torsion theoretic characterization of sheaves
which is exact. Restricting it to the full subcategory ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, we get a functor $\Gamma_{x}$ which is only left exact since the abelian structure of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is not inherited from that of $\mathcal{O}\operatorname{-Mod}$. Thus one can define $\mathrm{R}^{i}\Gamma_{x}(V)$, the $i$-th right derived functor of $\Gamma_{x}$, called the $i$-th sheaf cohomology group of $V$ at $x$. For details, please refer to [29, I.21.2].
@@
$\mathrm{R}^{i}\Gamma_{x}\cong\Gamma_{x}^{p}\circ(\mathrm{R}^{i+1}\mathcal{T}_{J}).$
@@
It has been proved in [8, Theorem 3.16] for the special case that $J$ is the atomic Grothendieck topology and $\mathcal{O}$ is a constant structure sheaf. The proof actually works in the general situation. ∎
@@
This is essentially a combination of [17, C2.3, Lemma 2.3.3] and [17, C2.3, Corollary 2.3.4], so we only give a sketch of the proof. The functor $\phi$ induces a restriction functor
@@ 5. Grothendieck topologies on artinian EI categories
The Serre quotient on the right hand side is still complicated, so we want to find a more explicit description for some special Grothendieck topologies $J$. When $J$ is a subcategory topology (or called rigid topology), this Serre quotient is equivalent to $\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$, where $\mathcal{O}_{\mathscr{D}}$ is the restriction of $\mathcal{O}$ to a full subcategory $\mathscr{D}$ of ${\mathscr{C}}$, [^4] Since $\mathcal{O}$ is a covariant functor from $\mathcal{C}$ to the category of associative rings, its shall be restricted to ${\mathscr{D}}$ rather than ${\mathscr{D}}^{{\mathrm{op}}}$. and hence becomes much more transparent. Furthermore, it has been pointed out in [2, IV.9] that all Grothendieck topologies on a finite Karoubi category are rigid, and this result also holds for artinian posets (see [22, Theorem 2.12]).
@@ 5.1. Definitions and the main result
When ${\mathscr{C}}$ is a directed category, one can construct a natural filtration on $\operatorname{Ob}({\mathscr{C}})$ as follows. Explicitly, let $O_{0}=\emptyset$, $O_{1}$ be the set of objects in $\operatorname{Ob}({\mathscr{C}})$ minimal with respect to $\leqslant$. For an ordinal $\chi$, let $O_{\chi}$ be the union of $O_{\chi-1}$ and the set of minimal elements in $\operatorname{Ob}({\mathscr{C}})\backslash O_{\chi-1}$ provided that $\chi$ is a successor ordinal, and let $O_{\chi}$ be the union of $O_{\lambda}$ for $\lambda<\chi$ if $\chi$ is a limit ordinal. In this way we obtain a chain $O_{1}\subseteq O_{2}\subseteq\cdots\subseteq\operatorname{Ob}({\mathscr{C}})$. Note that ${\mathscr{C}}$ is artinian if and only if there exists an ordinal $\zeta$ such that $O_{\zeta}=\operatorname{Ob}({\mathscr{C}})$. Dually, by taking maximal objects one gets another filtration $\{O^{\prime}_{\chi}\}$ on $\operatorname{Ob}({\mathscr{C}})$, and ${\mathscr{C}}$ is noetherian if and only if there exists an ordinal $\zeta$ such that $O^{\prime}_{\zeta}=\operatorname{Ob}({\mathscr{C}})$. For details, see [9, Subsection 3.1].
@@
Now we introduce the rigid topology and the main theorem of this section. Let ${\mathscr{C}}$ be a small skeletal category, and $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Recall from [17, C2.2] that $x\in\operatorname{Ob}({\mathscr{C}})$ is called $J$-irreducible if $J(x)=\{{\mathscr{C}}(x,-)\}$. For $y\in\operatorname{Ob}({\mathscr{C}})$, let $S_{y}$ be the sieve generated by the family of morphisms (in ${\mathscr{C}}$ rather than ${\mathscr{C}^{\mathrm{op}}}$) from $y$ to all $J$-irreducible objects. It can be explicitly described as follows
@@
A rigid Grothendieck topology $J$ is also called a subcategory topology with respect to the full subcategory consisting of $J$-irreducible objects. When $J$ is a rigid topology, $S_{y}$ is the unique minimal covering sieve in $J(y)$, so $J(y)=\{S\subseteq{\mathscr{C}}(y,-)\mid S\supseteq S_{y}\}$. For details, see [17, C2.2, Example 2.2.4(d), Remark 2.2.5, and Definition 2.2.18].
@@
Let ${\mathscr{C}}$ be a directed category such that ${\mathscr{C}}(x,x)$ is a finite set for every $x\in\operatorname{Ob}({\mathscr{C}})$. Then every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if ${\mathscr{C}}$ is a noetherian EI category.
@@ 5.2. The necessary direction
In this subsection we prove the only if direction of Theorem 5.7. Throughout this subsection let ${\mathscr{C}}$ be a directed category. As the starting point, we consider the special case that ${\mathscr{C}}$ has only one object, so we can identify ${\mathscr{C}^{\mathrm{op}}}$ with a finite monoid $M$, and correspondingly, a sieve on ${\mathscr{C}^{\mathrm{op}}}$ is nothing but a right ideal of $M$.
@@
A finite monoid $M$ has exactly two Grothendieck topologies if and only if $M$ is a finite group.
@@
Remark 5.10.
The assumption that $M$ is finite is essential to guarantee that $N$ is nonempty. Otherwise, let $M$ be the set of all surjections from an infinite set to itself. Clearly, $M$ is not a group, but every element in $M$ is right invertible. Thus the set $N$ is empty.
@@
Let ${\mathscr{C}}$ be a directed category such that every ${\mathscr{C}}(x,x)$ is finite. If every Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is rigid, then ${\mathscr{C}}$ is a noetherian EI category.
@@
We first show that ${\mathscr{C}}$ is an EI category; that is, for every $x\in\operatorname{Ob}({\mathscr{C}})$, the finite monoid $M_{x}={\mathscr{C}}(x,x)$ is a group. Let ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of objects $y$ with $y\geqslant x$. Clearly, ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$. By Lemma 5.11, every Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$ shall be rigid. Now let $\mathscr{E}$ be the full subcategory of ${\mathscr{D}}$ consisting of the single object $x$. Then $\mathscr{E}$ is a coideal of ${\mathscr{D}}$. By Lemma 5.12, every Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$ shall be rigid as well. By Lemma 5.9, the finite monoid $M_{x}^{{\mathrm{op}}}$ is a group, so is $M_{x}$. Therefore, ${\mathscr{C}}$ is an EI category.
@@
Remark 5.14.
The assumption that every ${\mathscr{C}}(x,x)$ is a finite set is only used in the proof of this proposition such that Lemma 5.9 can be applied.
@@ 5.3. Sufficient direction
(§)($\S$)( § ) each $\alpha\in S_{x}$ can be written as $\alpha=\beta\gamma$ with $\beta\in S_{y}$ and $\gamma\in T$.
@@
Now we check the claim (§)($\S$)( § ). Since $S\neq{\mathscr{C}}(x,-)$, by the consistent condition we can write $\alpha$ as a composite $\beta_{1}\gamma_{1}$ with $\gamma_{1}\in{\mathscr{C}}(x,x_{1})$ for a certain object $x_{1}\neq x$ and $\beta_{1}\in S_{x_{1}}$. We have two cases:
@@
Continuing the above procedure we obtain a sequence of objects $x<x_{1}<x_{2}<\ldots$ which must be finite since ${\mathscr{C}}$ is noetherian. Let $x_{n}$ be the last object, then either $S_{x_{n}}={\mathscr{C}}(x_{n},-)$ (and hence (§)($\S$)( § ) holds) or $x_{n}$ is maximal with respect to $\leqslant$. For the second case, we have $\beta_{n}\in S_{x_{n}}$ is an automorphism since ${\mathscr{C}}(x_{n},x_{n})$ is assumed to be a group, so $S_{x_{n}}={\mathscr{C}}(x_{n},-)$ as well. Consequently, the claim (§)($\S$)( § ) holds in both cases.
@@
(2): This is essentially [17, C2.2, Theorem 2.2.3]. For the convenience of the reader, we give a sketched proof. Given an $\mathcal{O}$-module $V$, its restriction to ${\mathscr{D}}$ is clearly an $\mathcal{O}_{{\mathscr{D}}}$-module. In this way we obtain an restriction functor $\operatorname{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$. It has a right adjoint given by the right Kan extension described as follows: given an $\mathcal{O}_{{\mathscr{D}}}$-module $W$ and an object $x\in\operatorname{Ob}({\mathscr{C}})$,
@@
where $\iota:{\mathscr{D}}\to{\mathscr{C}}$ is the inclusion functor, and the limit is taken over the comma category $(\iota\downarrow\mathrm{const}_{x})$. When $x\in\operatorname{Ob}({\mathscr{D}})$, it is not hard to see that $(\iota\downarrow\mathrm{const}_{x})$ is a coslice category whose objects are morphisms in ${\mathscr{D}}$ ending at $x$. In this situation, the identity morphism on $x$ is the initial object of the coslice category. It follows that $(\mathrm{Ran}_{\iota}W)_{x}\cong W_{x}$ for $x\in\operatorname{Ob}({\mathscr{D}})$. Consequently, $\operatorname{res}\circ\mathrm{Ran}_{\iota}$ is isomorphic to the identity functor on $\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}$. By [10, Proposition III.2.5], we conclude that
$\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}\simeq\mathcal{O}\operatorname{-Mod}/\ker(\operatorname{res})=\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J)\simeq{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})$
since the kernel of $\operatorname{res}$ is precisely $\mathcal{T}(J)$ by (1). ∎
@@
Note that the above result holds for any rigid Grothendieck topology $J$ on any small category ${\mathscr{C}}^{{\mathrm{op}}}$ by letting ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of $J$-irreducible objects; see [17, C2.2, Theorem 2.2.3] and [17, C2.2, Example 2.2.4(d)].
@@ 6. Grothendieck topologies on certain noetherian EI categories
[…]
@@
Let $J$ be an exceptional Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ corresponded to a function $\boldsymbol{d}$, and let ${\mathscr{D}}$ be the full subcategory consisting of objects $n$ such that $d_{n}=0$ or $d_{n}=\infty$. Then ${\mathscr{D}}^{{\mathrm{op}}}$ is a $J$-dense subcategory of ${\mathscr{C}}^{{\mathrm{op}}}$ (for a definition of $J$-dense subcategories, see [17, C2.2, Definition 2.2.1]). Therefore, applying the argument in the proof of Theorem 5.18, one has
@@
and $\mathcal{O}_{{\mathscr{D}}}$ is the restriction of $\mathcal{O}$ to ${\mathscr{D}}$. The full subcategory ${\mathscr{D}}$ is an artinian category whose objects can still be parameterized by $\mathbb{N}$, and $J_{{\mathscr{D}}}$ corresponds to the sequence (0,0,,0,,,)(0,0$\ldots$,0$\infty$,$\infty$,$\ldots$)( 0 , 0 , … , 0 , ∞ , ∞ , … ).
@@
One can deduce the above results from Lemma 5.11. Indeed, let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Then for each $n\in\mathbb{Z}$, the full subcategory ${\mathscr{C}}_{\geqslant n}$ of objects $m\geqslant n$ is an ideal of ${\mathscr{C}}$ and an EI category of type ${\mathbb{N}}$. These results follows via applying Propositions 6.3, 6.6 and 6.8 to the restriction of $J$ to ${\mathscr{C}}_{\geqslant n}$. When ${\mathscr{C}}$ is the poset $\mathbb{Z}$, they have been proved in [22, Proposition B.31].